OpenStudy (anonymous):

Help, and thank you in advance! (2/5n^9)^2 Came across this new problem not sure how to complete it.

OpenStudy (anonymous):

I'm supposed to simplify it. The options are, 10n^18 4/5n 4/10n^18 4/25n^18

OpenStudy (anonymous):

\[(\frac{ 2 }{ 5n^9 })^2 = \frac{ 2^2 }{ (5n^9)^2 } = \frac{ 4 }{ 25n^{18} }\] exponents multiplied since \[(x^{a})^{b} = x^{ab}\]

OpenStudy (anonymous):

Ok thank you very much I understand :)